Show work please

You purchase one (1) call option with strike price 50 for $ 9 and write three (3) call options with strike 60 for $ 3.

1) Draw the payoff and profit table for this strategy at maturity.

2) When do you break-even (profit=0) at maturity?

3) What are your anticipations about the stock at maturity (when do you make money)?

4) Assume that you may purchase calls with strike price 70 for $ 1. How many options would you trade to prevent unbounded losses at maturity? What would be the maximum extent of your losses after the purchase?
Answer #1

Premium for purchase = -9

Premium for writing off 3 call = 3*3 =9

So total premium = -9+9 =0

Example if share price is 55

then call price of $50 Strike price will be = Share price- Strike price = 55-50 =5

then call price of $60 Strike price will be = Share price- Strike price or zero if Share price < Srike price  =0

SO total payoff = Value of $50 call price +value of $60 call price + Premium = 5+0+0 =$5

Similar way you can construct table as shown below :

a)Payoff table without formula

50 60
Expected share price Payoff from 1 call option Payoff from 3 call option SP $60 Premium Total profit or loss
35 0 0 0 0
40 0 0 0 0
45 0 0 0 0
50 0 0 0 0
55 5 0 0 5
60 10 0 0 10
65 15 -15 0 0
70 20 -30 0 -10
75 25 -45 0 -20
80 30 -60 0 -30

b) Payoff function withforomula

50 60
Expected share price Payoff from 1 call option Payoff from 3 call option SP $60 Premium Total profit or loss
35 =IF(A15>$B$13,A15-$B$13,0) =3*IF(A15<$C$13,0,-A15+$C$13) =-9+3*3 =D15+C15+B15
40 =IF(A16>$B$13,A16-$B$13,0) =3*IF(A16<$C$13,0,-A16+$C$13) =-9+3*3 =D16+C16+B16
45 =IF(A17>$B$13,A17-$B$13,0) =3*IF(A17<$C$13,0,-A17+$C$13) =-9+3*3 =D17+C17+B17
50 =IF(A18>$B$13,A18-$B$13,0) =3*IF(A18<$C$13,0,-A18+$C$13) =-9+3*3 =D18+C18+B18
55 =IF(A19>$B$13,A19-$B$13,0) =3*IF(A19<$C$13,0,-A19+$C$13) =-9+3*3 =D19+C19+B19
60 =IF(A20>$B$13,A20-$B$13,0) =3*IF(A20<$C$13,0,-A20+$C$13) =-9+3*3 =D20+C20+B20
65 =IF(A21>$B$13,A21-$B$13,0) =3*IF(A21<$C$13,0,-A21+$C$13) =-9+3*3 =D21+C21+B21
70 =IF(A22>$B$13,A22-$B$13,0) =3*IF(A22<$C$13,0,-A22+$C$13) =-9+3*3 =D22+C22+B22
75 =IF(A23>$B$13,A23-$B$13,0) =3*IF(A23<$C$13,0,-A23+$C$13) =-9+3*3 =D23+C23+B23
80 =IF(A24>$B$13,A24-$B$13,0) =3*IF(A24<$C$13,0,-A24+$C$13) =-9+3*3 =D24+C24+B24

Note: Incase of any doubt, please do comment. I will get back to you. Kindly post rest of the questions sapratly since as per policy I can not anwer more than 1 question. Thanks!!